OpenStudy (anonymous):

The prism shown below will be sliced by a plane perpendicular to the base along the dotted line AB. Points A and B are mid points of the equal edges

OpenStudy (anonymous):

What will be the approximate area of the shaded cross section? 57 cm2 112 cm2 80 cm2 70 cm2

OpenStudy (allank):

I'd go for 57 cm2. Here is my logic: The line AB seems to be approximately the hypotenuse of a right triangle with legs of length 5cm and 5cm. Thus, applying Pythagorean theorem, AB = sqrt(5^2+5^2) = sqrt(50) Therefore the shaded area = AB*8 = sqrt(50)*8 = 56.67 which is approximately 57 cm2